OpenStudy (lilmike234):

let log of PN6; log of MN=8; what is the value

OpenStudy (lilmike234):

No P/N and M/N

OpenStudy (dumbcow):

ok \[\ln (p/n) = 6, \ln (m/n) = 8\]

OpenStudy (lilmike234):

What's the relationship ? Is it M=100P

OpenStudy (dumbcow):

using the log property ..... log a - log b = log(a/b) \[\ln p - \ln n = 6 \rightarrow \ln n = \ln p - 6\] \[\ln m - \ln n = 8\] using substitution \[\ln m - (\ln p - 6) = 8\] \[\ln m - \ln p = 2\] \[\ln (m/p) = 2\] .... assuming log base 10, then yes m/p = 100 or m = 100p